OpenStudy (anonymous):

What are the solutions to the equation x2 = 81?

OpenStudy (anonymous):

Square root to both sides . \[\huge \sqrt{x^2}=\sqrt{81}\]

OpenStudy (anonymous):

When you find the square root, you get both a positive and negative answer. Therefore, x^2 = 81 x = √81 x = ±9

OpenStudy (anonymous):

x=+-9

OpenStudy (anonymous):

When you find the square root, you get both a positive and negative answer. Therefore, x^2 = 81 x = √81 x = ±9

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome! :)

OpenStudy (anonymous):

x = 40.5 is my guess

OpenStudy (lgbasallote):

@Sara13 dont forget to give a medal (by pressing best answer) to the people who gave you the answer that satisfied you the most ;) it's a sign of appreciation and respect ehre

OpenStudy (anonymous):

the other side is 32